An impatient motorist considers speeding as he travels between two cities. If the trip normally takes 0.8h at an average speed of 100km/h how much time will be saved if he exceeds the limit by 10km/h? Answer in minutes.
1
Expert's answer
2019-09-24T09:08:02-0400

The distance:


"d=vt=(100)(0.8)=80\ km"

New time:


"T=\frac{d}{V}=\frac{80}{100+10}=\frac{8}{11}h"


"t-T=0.8-\frac{8}{11}=0.073 \ h=4.4\ min"